express the following statement as a formula with the value of the constant of proportionality determined with the given conditions: w varies directly as x and inversely as the square of y. if x=15 and y=5, then w=36.
Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND TRY...USE W FOR R...X FOR L....AND...Y FOR D.
the elctrical resistance R of a wire varies directly as its length L and inversely as the sqaure of its diameter. A wre 20 meters long and 0.6 centimeters in diameter made from a certain alloy has a resistance of 36 ohms. what is the resistance of a pice of wire 60 meters long and 1.2 centimeters in diameter made from the same material?
R = KL/D^2....WHERE K IS A CONSTANT
36=K*20/0.6^2
K=36*0.36/20
R=K*60/1.2^2=[36*0.36/20][60/1.44]=[36*0.36*60/(20*1.44)]=27 OHMS